# Unveiling the Learning Mind of Language Models: A Cognitive Framework and Empirical Study

### v1

@@ 7 LearnArena: A Benchmark Suite for General Learning Ability
8 Conclusion
To address the lack of analyzing of learning ability in large language models, we propose a cognitively grounded framework that decomposes general learning into three key dimensions: learning from instructor, learning from concept, and learning from experience. This decomposition enables a principled analysis of how LLMs acquire, internalize, and apply new knowledge. Our experiments reveal that interaction enhances instruction-based learning, conceptual understanding scales with model size, and LLMs struggle in many-shot scenarios due to long-context limitations. Motivated by these findings, we introduce LearnArena, a unified benchmark designed to evaluate general learning abilities across cognitively meaningful settings.
Impact Statements
This work provides a novel, cognitively grounded framework for evaluating the general learning ability of large language models (LLMs), addressing a critical yet underexplored dimension of LLM capabilities. By decomposing learning into instructor-based, conceptual, and experiential dimensions, it offers a structured lens for diagnosing model behavior and guiding model development. The proposed benchmark facilitates realistic and principled evaluation of adaptive learning, benefiting the design of more robust, human-like AI systems in dynamic and open-ended environments. This work has broad implications for AI alignment, continual learning, and educational AI applications.
Appendix A Experimental Setup
A.1 Implementation Details
All experiments were conducted on a high-performance computing server equipped with an AMD EPYC 7V12 64-Core Processor, 512 GB of system memory, and 96 logical CPUs. The system also includes eight NVIDIA A100 GPUs, each with 80 GB of memory. Our training script is adapted from the example provided by LlamaFactory (Zheng et al., 2024) [^1] https://github.com/hiyouga/LLaMA-Factory/blob/main/examples/train_full/llama3_full_sft_ds3.yaml Full configuration details are provided in Appendix E. The references to Llama3.1-8B, Qwen2.5-7B, Mistral-7B, Phi-3-mini, Qwen2.5-1.5B, Qwen2.5-14B, Qwen2.5-32B, Qwen2.5-72B, Mistral-8B, Qwen3-8B, and Qwen3-14B in the main text refer to the specific models: Llama-3.1-8B-Instruct, Qwen2.5-7B-Instruct, Mistral-7B-Instruct-v0.3, Phi-3-mini-128k-instruct, Qwen2.5-1.5B-Instruct, Qwen2.5-14B-Instruct, Qwen2.5-32B-Instruct, Qwen2.5-72B-Instruct, Mistral-8B-Instruct-2410, Qwen3-8B, and Qwen3-14B, respectively.
A.2 Evaluation Datasets
SVAMP
(Patel et al., 2021) SVAMP tasks are designed to provide a robust benchmark for evaluating the mathematical reasoning capabilities of language models on Math Word Problems (MWPs). Each task presents a short natural language narrative that describes a real-world scenario and poses a quantitative question. Previous benchmarks, such as ASDiv-A and MAWPS, were shown to permit high model performance even when critical question components were removed or word-order was disregarded, revealing the presence of shallow heuristics in existing approaches. SVAMP addresses these deficiencies by introducing examples that require models to demonstrate sensitivity to the question, robust reasoning, and invariance to structural alterations.
MATH
(Hendrycks et al., 2021b) MATH tasks are constructed from a dataset of 12,500 challenging competition-level mathematics problems, specifically designed to evaluate the mathematical problem-solving capabilities of language models. Each problem is accompanied by a detailed, step-by-step solution, enabling both the assessment of answer derivations and the development of explanation generation. In addition to the main benchmark, a large auxiliary pretraining dataset is provided to help teach models the fundamentals of mathematics. Despite advancements in large-scale Transformer models, accuracy on the MATH benchmark remains relatively low, highlighting the limitations of current approaches and the need for new algorithmic innovations in mathematical reasoning.
NumGLUE
(Mishra et al., 2022) NumGLUE tasks are derived from a comprehensive multi-task benchmark designed to evaluate the numerical reasoning abilities of AI systems across eight distinct task types, each requiring some form of arithmetic understanding. The benchmark includes tasks involving commonsense and domain-specific arithmetic reasoning, quantitative comparison, fill-in-the-blank questions, reading comprehension with explicit and implicit numerical reasoning, quantitative natural language inference, and arithmetic word problems. By assessing model performance across these diverse task formats, NumGLUE provides a rigorous evaluation of both simple and complex numerical reasoning skills without leveraging explicit task-type labels during training.
SimulEq
(Koncel-Kedziorski et al., 2016) SimulEq tasks are based on a dataset comprising 514 test samples centered on equation solving, with a particular focus on algebraic manipulation and logical reasoning. The dataset is designed to evaluate models’ proficiency in formulating and solving algebraic equations through the application of structured reasoning processes.
AQuA
(Ling et al., 2017) AQuA tasks are derived from the AQUA-RAT dataset, which comprises approximately 100,000 algebraic word problems accompanied by natural language rationales. Each problem includes a natural language question, five answer options (A–E) with a single correct choice, and a detailed rationale describing the solution process. This dataset enables evaluation of both answer selection and reasoning transparency in mathematical problem solving.
SAT
(Zhong et al., 2023) SAT tasks are sourced from the AGIEval benchmark, which is designed to evaluate the general abilities of foundation models using authentic, human-centric standardized exams. The SAT component consists of mathematics and verbal reasoning problems that mirror those found in real-world college entrance assessments. By incorporating these tasks, AGIEval provides a robust framework for assessing models’ understanding, reasoning, and calculation skills at a human-competitive level. Recent results show that state-of-the-art models, such as GPT-4, can achieve or surpass average human performance on SAT tasks, highlighting both the progress and current limitations of large language models in solving complex, human-oriented problems.
MMLU-Math
(Hendrycks et al., 2020a) MMLU-Math tasks are part of a comprehensive multitask benchmark designed to evaluate a model’s accuracy across a wide range of academic subjects, with a specific focus on elementary and advanced mathematics. The benchmark requires models to demonstrate both broad world knowledge and sophisticated problem-solving skills. While recent large language models show improvements over random-chance accuracy, substantial gaps remain in achieving expert-level performance, particularly in mathematics. MMLU-Math enables detailed analysis of a model’s mathematical proficiency and helps to identify critical areas requiring further advancement.
MagpieMath
(Xu et al., 2024a) MagpieMath is a large-scale dataset generated via the Magpie pipeline, which synthesizes high-quality mathematical instruction data through auto-regressive sampling from Qwen2.5-Math. Unlike prior instruction tuning datasets that depend on seed prompts or handcrafted templates, Magpie employs a pre-query-based sampling approach to elicit diverse and pedagogically rich instruction data directly from aligned LLMs. The resulting dataset spans a wide range of mathematical skills, including arithmetic, algebra, and multi-step reasoning, and serves as a robust resource for evaluating instruction-based learning in mathematical domains.
Poker
(Guertler et al., 2025) is a strategic two-player Texas Hold’em game where players compete by betting chips based on their hole cards and community cards. Each player aims to win chips by either having the best hand at showdown or making their opponent fold. The game follows standard Texas Hold’em rules with a fixed number of rounds and betting structures.
Ultimate Tic Tac Toe
(Guertler et al., 2025) is a strategic two-player game that combines the classic Tic Tac Toe with an added layer of complexity. Players aim to win three micro boards in a row (horizontally, vertically, or diagonally) on the macro board, which tracks the outcomes of individual micro boards. Each move influences the opponent’s next playable micro board, creating dynamic and strategic gameplay. This environment implements the full rules of Ultimate Tic Tac Toe, including valid move enforcement, micro board and macro board win detection, and a clear rendering of the board state for agent-based gameplay and experimentation.
Checkers
(Guertler et al., 2025) (also known as Draughts) is a turn-based game played on an 8×8 board. Each player controls a set of pieces placed on alternating dark squares. Pieces move diagonally forward and can capture opponent pieces by jumping over them. A piece that reaches the opposite side of the board becomes a "King," gaining the ability to move backward as well. The game ends when one player’s pieces are all captured or when a player has no valid moves left.
Stratego
(Guertler et al., 2025) is a two-player strategy board game where players aim to capture their opponent’s Flag or eliminate all of their movable pieces. Each player strategically places their army, consisting of pieces of varying ranks, on a 10x10 grid, with special rules for movement and battle. Bombs act as immovable traps, while Scouts can move multiple spaces in a straight line. The game involves hidden information, where a player’s pieces are unknown to their opponent until revealed in battle. This environment simulates the full rules of Stratego, including strategic placement, battle resolution, and dynamic board rendering, providing a robust setup for agent-based gameplay and experimentation.
TicTacToe
(Guertler et al., 2025) (also known as Noughts and Crosses) is a classic two-player game played on a 3×3 grid. Players take turns marking cells with their symbol (’X’ or ’O’), with the objective of placing three of their marks in a horizontal, vertical, or diagonal row. This implementation provides a complete game environment with move validation, win detection, and draw conditions.
Truth and Deception
(Guertler et al., 2025) is a two-player social deduction game where one player (the Deceiver) attempts to mislead the other player (the Guesser) about which of two facts is correct. The players engage in a conversation of limited turns, during which the Deceiver strategically shares information to encourage the Guesser to choose the incorrect fact. At the end of the conversation, the Guesser must determine which fact is true. The game tests persuasion, critical thinking, fact verification, and strategic communication in a structured environment with clear objectives for both roles.
Spelling Bee
(Guertler et al., 2025) is a two-player word game where players take turns creating valid English words using only a limited set of allowed letters. The challenge increases as the game progresses because each new word must be at least as long as the previous word. Players must balance vocabulary knowledge with strategic thinking to outmaneuver their opponent by forcing them into increasingly difficult positions. This implementation features frequency-weighted letter selection to ensure balanced and playable letter sets.
Spite and Malice
(Guertler et al., 2025) is a two-player competitive card game that combines elements of solitaire and strategic play. Each player has their own payoff pile that they aim to deplete first to win. Players can play cards to shared center piles in ascending sequence (Ace to Queen), with Kings serving as wild cards. The game involves careful resource management, opportunistic card placement, and strategic blocking to prevent your opponent from emptying their payoff pile. This implementation features a complete deck management system, discard piles, and a hand limit of five cards.
Tak
(Guertler et al., 2025) is a two-player abstract strategy board game. Players aim to build a continuous road connecting two opposite edges of the board while blocking their opponent’s attempts. The game involves placing and moving pieces of various types—Flat Stones, Standing Stones (Walls), and Capstones—each with unique abilities. This environment simulates the full rules of Tak, including flexible board sizes, road-building mechanics, and dynamic board rendering, offering a rich platform for agent-based gameplay and experimentation.
Word Chains
(Guertler et al., 2025) is a turn-based word game where two players take turns providing valid English words that start with the last letter of the previous word. The game enforces rules such as no word repetition and requires valid English words. Players lose if they fail to provide a valid word.
LogicGrid
(Mandell, 1986) LogicGrid tasks are constructed as logic puzzles that require participants to perform grid-based deduction and symbolic reasoning. These tasks are specifically designed to evaluate the systematic application of abstract rules within constrained environments. All LogicGrid tasks presented in this work are newly developed adaptations based on established principles from classic puzzle collections.
NaLogic
(Dudeney, 1907) NaLogic tasks are narrative-driven logic puzzles that require solvers to extract and apply logical constraints embedded within rich, contextualized storytelling. These tasks are directly inspired by literary puzzle formats, and they are specifically designed to assess the integration of logical reasoning within narrative comprehension. All NaLogic tasks included in this work are systematically derived from established approaches found in classic puzzle literature.
Plan
(Vallati et al., 2015) Plan tasks are adapted from classical scenarios formulated within the Planning Domain Definition Language (PDDL), including representative domains such as Gripper, Barman, Blocksworld, and Tyreworld. In these tasks, symbolic states are systematically translated into natural language descriptions to enable seamless interaction with large language models. Progress is quantitatively tracked by computing state-goal matching scores, thus providing a robust measure of task completion and logical reasoning within dynamic planning environments.
AlfWorld
(Shridhar et al., 2021) AlfWorld tasks require agents to perform everyday household activities that necessitate both exploration and commonsense reasoning. AlfWorld leverages interactive TextWorld environments that are closely aligned with the embodied scenarios found in the ALFRED dataset. This alignment enables agents to develop and reason about high-level policies within an abstract, text-based space, prior to executing embodied tasks that demand low-level actuation in physical or simulated environments.
ScienceWorld
(Wang et al., 2022a) ScienceWorld tasks require agents to comprehend and execute a variety of scientific procedures, including activities such as measuring physical properties and manipulating laboratory equipment. Each task is accompanied by carefully reannotated subgoals designed to more accurately capture incremental progress, thereby enabling a granular evaluation of scientific reasoning and procedural understanding.
BabyAI
(Chevalier-Boisvert et al., 2019) BabyAI tasks require agents to navigate and interact within grid-world environments, leveraging newly introduced textual action representations and re-annotated subgoals. These modifications are designed to facilitate more nuanced agent-environment interaction and provide finer-grained evaluation of task completion in language-conditioned navigation settings.
AlpacaEval
(Dubois et al., 2024) AlpacaEval tasks consist of a diverse set of 805 instructions, systematically aggregated from multiple benchmark sources. Specifically, the collection includes 252 instructions from the self-instruct test set(Wang et al., 2022b), 188 from the Open Assistant test set (OASST), 129 from the Anthropic Helpful Test Set (Zhou et al., 2023), 80 from the Vicuna test set (Chiang et al., 2023), and 156 from the Koala test set (Vu et al., 2023). This comprehensive assembly is designed to support the broad evaluation of the instruction-following capabilities across a wide range of language models.
Vicuna
(Chiang et al., 2023) Vicuna tasks consist of 80 test instructions, systematically organized into eight distinct categories: Fermi problems, common sense reasoning, role-play scenarios, coding, mathematics, writing tasks, counterfactuals, knowledge assessment, and generic questions. This categorical organization is designed to facilitate a thorough evaluation of multiple dimensions of chatbot performance. Prior research suggests that the Vicuna dataset predominantly features instructions of comparatively lower difficulty and complexity(Xu et al., 2023). In this study, the Vicuna test set is used to specifically assess the performance of large language models across these various types of instruction.
Self-Instruct
(Wang et al., 2022b) Self-Instruct tasks consist of 252 human-authored test instructions, each paired with a carefully crafted reference output. This test set is meticulously curated to represent the real-world applicability of instruction follower models, encompassing a wide range of domains such as email composition, social media engagement, productivity software, and coding tasks. The instructions exhibit substantial diversity in both style and format, including various task lengths and a spectrum of input/output types such as bullet points, tables, code fragments, and mathematical equations. In this study, the Self-Instruct test set is utilized to rigorously evaluate the model’s proficiency in adhering to precise instructions across heterogeneous domains.
Wizardlm
(Xu et al., 2023) Wizardlm tasks are derived from a large-scale instruction dataset initially seeded with 52,000 instructional examples from Alpaca, subsequently expanded through $M=4$ evolutionary cycles to produce 250,000 unique instructions. In each cycle, for every instruction, six new prompts are generated—five via in-depth evolution and one through in-breadth evolution—of which one prompt is randomly selected. Responses are then generated using ChatGPT, culminating in $52,000\times 4\times 3=624,000$ instruction-response pairs. The training subset utilized for the Evol-Instruct dataset comprises 70,000 of these instructions. The test set, consisting of 218 instructions, is curated from diverse sources, including open-source projects and online forums, and encompasses 29 distinct skill areas identified from real-world human tasks.
ARC
(Clark et al., 2018) ARC tasks are derived from a dataset comprising 7,787 English-language science exam questions sourced from a variety of educational materials, including licensed content provided by a research partner affiliated with AI2. Each question is presented in a multiple-choice format, typically with four answer options, and spans a range of grade levels. The dataset is partitioned into a Challenge Set containing 2,590 “hard” questions—defined as those that cannot be correctly answered by either retrieval-based or co-occurrence-based baseline methods—and an Easy Set containing 5,197 questions of comparatively lower difficulty. This structure supports comprehensive evaluation of reasoning and scientific understanding across different levels of complexity.
HellaSwag
(Zellers et al., 2019) HellaSwag tasks are derived from a challenging benchmark for commonsense natural language inference, wherein models must select the most plausible continuation for a given event description. For example, given an event such as “A woman sits at a piano,” the model must choose the most likely follow-up, such as “She sets her fingers on the keys.” Despite the introduction of advanced pretrained models like BERT, which achieved near human-level performance on previous tasks, HellaSwag demonstrates that even state-of-the-art models still struggle with robust commonsense reasoning. The dataset is constructed using Adversarial Filtering, a methodology where a series of discriminators iteratively select adversarially challenging, machine-generated distractors. While these questions are trivial for humans (accuracy $>$95%), state-of-the-art models frequently misclassify them (accuracy $<$48%).
WinoGrande
(Sakaguchi et al., 2019) WinoGrande tasks are derived from a dataset comprising 44,000 problems, inspired by the Winograd Schema Challenge (Kocijan et al., 2023), but specifically adjusted to enhance both scale and robustness against dataset-specific biases. Each task is presented as a fill-in-the-blank question with binary options, requiring the participant to select the correct answer based on contextual commonsense reasoning.
MMLU
(Hendrycks et al., 2021a) MMLU tasks are derived from the benchmark introduced by Hendrycks et al., evaluating a model’s ability to answer multiple-choice questions across a broad range of academic subjects. In the original format, each question presents the answer choices within the context, and the model is required to produce the answer letter (e.g., A, B, C, D) as the continuation. The mmlu-continuation variant adopts a cloze-style format, omitting the answer choices in the context and requiring the full answer choice in the continuation. The mmlu-generation variant is similar to the original, but the model is tasked with generating the correct answer letter directly. These variants collectively enable a nuanced evaluation of language models’ reasoning and generation capabilities across diverse knowledge domains.
GSM8k
(Cobbe et al., 2021) GSM8k tasks are drawn from a dataset of 8,500 high-quality, linguistically diverse grade school math word problems, designed to evaluate multi-step mathematical reasoning in language models. Despite the conceptual simplicity of the problem distribution, state-of-the-art transformer-based models consistently struggle to achieve high performance on this benchmark. The GSM8k dataset thus provides a rigorous diagnostic tool for analyzing the limitations of current models and supporting further research in mathematical reasoning.
Intlaw
(Oppenheim, 1920, 1921) Intlaw tasks are derived from L. Oppenheim’s two-volume International Law: A Treatise, and are designed to assess understanding of both foundational and applied principles in international law. Volume 1 emphasizes legal theory, sources of law, and the concept of state sovereignty, while Volume 2 explores topics such as conflict, neutrality, and mechanisms for peaceful dispute resolution. This dataset enables evaluation of models’ capabilities in legal reasoning, state interactions, and comprehension of diplomatic and legal frameworks.
Willidea
(Schopenhauer, 1883, 1886, 1888) Willidea tasks are constructed from all three volumes of Arthur Schopenhauer’s The World as Will and Idea, focusing on the evaluation of philosophical reasoning and metaphysical analysis. The dataset challenges models to interpret complex philosophical arguments, engage with Kantian critique, and reason across abstract conceptual structures related to perception, will, and reality. Through these tasks, the ability of language models to comprehend and analyze dense philosophical discourse is rigorously assessed.
Ancsoc
(Morgan, 1877) Ancsoc tasks are derived from Lewis H. Morgan’s Ancient Society and are designed to assess a model’s understanding of the evolution of human civilization through the stages of savagery, barbarism, and civilization. The dataset emphasizes analysis of social structures, cultural development, and key concepts in anthropological theory. Through these tasks, models are evaluated on their capacity for historical analysis and cross-cultural reasoning.
Poliec
(Roscher, 1878, 1882) Poliec tasks are based on Wilhelm Roscher’s two-volume Principles of Political Economy and focus on the historical development of economic systems, classification of income, and the influence of human behavior on economic theory. The dataset is designed to assess models’ abilities to reason through complex economic frameworks, drawing on both historical context and theoretical foundations.
A.3 Details of Learning from the Instructor
We evaluate instruction-based learning ability across five distinct domains: Mathematics, Law, Philosophy, Anthropology, and Economy. Each domain is selected to reflect different facets of instruction-following behavior, ranging from structured problem-solving to conceptual understanding. During training, we additionally use the MagpieMath dataset (Xu et al., 2024a), which provides large-scale, high-quality mathematical instruction data synthesized from Qwen2.5-Math. It is a subset of the dataset available at this link  [^2] https://huggingface.co/datasets/Magpie-Align/Magpie-Qwen2.5-Math-Pro-300K-v0.1. The dataset is also accessible through our anonymous link  [^3] https://anonymous.4open.science/r/LRB-1FEA. The prompt can be found in Appendix  H. We used eight representative test benchmarks: GSM8K (Cobbe et al., 2021), SVAMP (Patel et al., 2021), MATH (Hendrycks et al., 2021b), NumGLUE (Mishra et al., 2022), SimulEq (Koncel-Kedziorski et al., 2016), AQuA (Ling et al., 2017), SAT (Zhong et al., 2023), and MMLU-Math (Hendrycks et al., 2020a). These datasets cover a range of mathematical competencies from grade-school arithmetic to advanced problem-solving, including multi-step reasoning, equation solving, and standardized testing skills. In the Law domain, we use the Intlaw dataset (Oppenheim, 1920, 1921). It covers core topics such as sovereignty, conflict, and dispute resolution, and evaluates models’ legal reasoning and understanding of international legal systems. In the Philosophy domain, we use the Willidea dataset (Schopenhauer, 1883, 1886, 1888), which focuses on metaphysics, perception, and will, testing models’ ability to interpret abstract arguments and philosophical reasoning. In the Anthropology domain, we use the Ancsoc dataset (Morgan, 1877), which evaluates understanding of social evolution, cultural systems, and anthropological theory. In the Economy domain, we use the Poliec dataset (Roscher, 1878, 1882), which emphasizes historical and theoretical reasoning in value, income, and economic behavior. Across these four non-math domains, question formats included fill-in-the-blank (fill), judgment (judge), multiple-choice (multi-choice), open-ended (open), and single-choice (single-choice). We provide representative cases of each question format in Appendix F. Further dataset details are provided in Appendix A.2
A.4 Details of Learning from the Concept
We evaluate models’ ability to learn from structured conceptual information across a variety of tasks that require rule interpretation, logical reasoning, and symbolic planning. Our experiments focus on two complementary settings: multi-agent competitive environments and logic-based planning tasks. In each case, we compare model performance with and without access to high-level conceptual summaries that describe task-specific rules, goals, or strategies. These summaries are generated by a stronger teacher model and provided as auxiliary context to the learner model. We first study conceptual learning in strategic settings using the TextArena framework (Guertler et al., 2025), a collection of multi-agent environments characterized by symbolic rules and multi-turn dynamics. Each game involves two players, Player-0 and Player-1. We fix Player-0 as Qwen2.5-32B and vary Player-1 across four model scales: Qwen2.5-1.5B, 7B, 14B, and 32B. Prior to gameplay, Player-1 either receives no guidance (without Concept) or is given a natural language description of the game rules and strategies (with Concept) generated by Qwen2.5-32B. For each environment and model combination, we run 20 matches and report the average win rate of Player-1. The evaluation environments include Checkers (CH), Poker (PK), Stratego (ST), Tic Tac Toe (TT), Truth and Deception (TD), and Ultimate Tic Tac Toe (UTT). Performance is measured as the win rate of Player-1 against the fixed opponent. We then evaluate rule-based generalization across a set of logic and planning tasks. The LogicGrid and NaLogic tasks consist of structured and narrative logic puzzles (Mandell, 1986; Dudeney, 1907), requiring deduction under symbolic constraints. Plan (Vallati et al., 2015) adapts classical PDDL-style planning domains (e.g., Gripper, Barman) into natural language to assess symbolic action reasoning. Additionally, we include AlfWorld (Shridhar et al., 2021), ScienceWorld (Wang et al., 2022a), and BabyAI (Chevalier-Boisvert et al., 2019) tasks, which involve procedural and goal-oriented behaviors requiring commonsense reasoning and spatial understanding. For the concept-injected setting, we use Qwen2.5-72B to generate static conceptual summaries tailored to each task. These summaries are provided as additional input to the learner model during evaluation. Further dataset details are provided in Appendix A.2
A.5 Details of Learning from the Experience
We investigate how language models adapt their behavior based on prior experiences. Our experiments are designed to capture different aspects of experience-based learning, ranging from competitive multi-turn interactions to static in-context supervision. In all settings, models are exposed to past interaction traces or curated examples, which serve as proxies for experiential knowledge. We evaluate two complementary scenarios: (1) interactive adaptation through self-selected experience in multi-agent games, and (2) passive conditioning on externally provided in-context examples. These two settings align conceptually with classical learning paradigms: the first resembles on-policy learning, where the model actively explores and adapts through its own interactions; the second mirrors off-policy learning, where the model passively leverages prior trajectories without influencing their generation. Together, they allow us to explore whether LLMs can extract useful patterns from previous data. We adopt the same multi-agent environment setup described in Section 5.1, using TextArena (Guertler et al., 2025), a suite of competitive games involving symbolic rules and multi-turn dynamics. The results are shown in Figure 5. In this setting, "with experience" refers to a scenario where, during the $k$-th game, the model selects three prior experiences from rounds 0 to $k{-}1$ (or all available ones if fewer than three exist). These samples are used as in-context learning (ICL) examples, allowing the model to draw on its past trajectories for real-time adaptation. Performance is measured by the average win rate of Player-1 across 20 independent matches. This setup captures whether models can benefit from self-curated experiential memory in competitive decision-making. To further analyze how accumulated experience affects model behavior, we treat in-context examples as off-policy trajectories that encode supervision signals. This setup parallels off-policy learning: unlike competitive gameplay in Section 6.1, where models adapt through active interaction (on-policy), here they condition passively on externally provided experience examples. These exemplars are not generated by the current model but are instead drawn from prior runs or curated datasets. We use MagpieMath (Xu et al., 2024a), introduced in Section4, as the source of experience for in-context learning, with both the ICL examples and test set drawn from this dataset. All in-context experiments are repeated 20 times using different random seeds. We report the mean and standard deviation for each result. Further dataset details are provided in Appendix A.2
A.6 Details of LearnArena
[…]

### v3

@@ A.6 Details of LearnArena
Appendix B Opponent Robustness and Extended Evaluation
LearnArena in the main text fixes Player-0 to Qwen2.5-32B. To assess robustness to opponent choice and reduce the risk that models overfit to a single opponent’s idiosyncrasies, we repeated the evaluation with three alternative Player-0 models—Phi-4, GPT-4o-mini, and GPT-4o. For each fixed Player-0, we measured Player-1 performance on four environments (Checkers, Stratego, SpellingBee, SpiteAndMalice) under the same protocol as the main paper, averaged win rates across the four environments to obtain a per–Player-1 score, ranked Player-1 models by this score, and computed the Spearman rank correlation between this ranking and the baseline ranking obtained with Player-0 = Qwen2.5-32B.
Table 3: Player-0 = Phi-4. Player-1 win rates on four environments.
Player-1 CH ST SB SM
gpt-4o 1.00 0.90 0.65 1.00
gpt-4o-mini 0.65 0.88 0.86 0.72
llama-3.1-8b 0.65 0.70 0.50 0.75
mistral-8b 0.40 0.75 0.28 0.69
qwen2.5-7b 0.70 0.75 0.35 0.35
qwen3-8b 0.75 0.65 0.50 0.70
Table 4: Player-0 = GPT-4o-mini. Player-1 win rates on four environments.
Player-1 CH ST SB SM
gpt-4o 0.80 1.00 0.65 1.00
gpt-4o-mini 0.50 0.55 0.45 0.55
llama-3.1-8b 0.25 0.85 0.10 0.70
mistral-8b 0.15 0.85 0.20 0.45
qwen2.5-7b 0.45 0.70 0.25 0.35
qwen3-8b 0.40 0.95 0.55 0.50
Table 5:
Across these alternative opponents, the rank-order agreement with the baseline is high. The Spearman correlations with the Qwen2.5-32B–based ranking are $\rho_{\text{Phi-4}}$ = 1.000,  $\rho_{\text{GPT-4o-mini}}$ = 0.829,  $\rho_{\text{GPT-4o}}$ = 0.714. These values indicate that relative model rankings are largely preserved across substantially different opponents. In practice, stronger opponents compress absolute win rates yet retain the broad ordering of Player-1 models, which is useful for difficulty calibration while maintaining comparative validity. Taken together, these results suggest that the main conclusions of the benchmark are not overly sensitive to the specific choice of Player-0. While average win rate remains a clear and discriminative summary of success in dynamic environments, it is a coarse measure of learning. To provide a more granular view of learning dynamics in future versions of LearnArena (LearnArena), we will augment reporting with process-level signals that capture how models adapt, follow guidance, and reuse experience. Specifically, we will measure adaptation speed through per-round improvement and rounds-to-threshold; instruction adherence through agreement between instructor feedback and subsequent actions under automatic alignment checks; concept utilization through rule compliance and strategy-template coverage when conceptual summaries are provided; experience reuse through the recurrence of salient patterns from selected prior trajectories in current decisions; and stability and robustness through variance across seeds and sensitivity to controlled noise injected into instructions, concepts, and experiences.
Appendix C Many-Shot Behavior and Long-Context Models
This section examines whether degradation in many-shot settings is merely a consequence of poor long-context capability. In the many-shot experiments reported in Section 6.2, we evaluate Qwen2.5-7B/14B-Instruct-1M, which support up to 1M-token windows and are explicitly designed for long-context usage. Across these models, performance follows a clear non-monotonic trend: accuracy improves as in-context examples increase, peaks at a small number of shots (3–9 depending on model size), and then declines as the prompt grows longer. This indicates that, even with extended context windows, current models remain effective few-shot learners while many-shot generalization remains challenging. To further validate this pattern with a frontier model under tractable cost, we additionally evaluate GPT-4o-mini (128k context) following the protocol in Section 6.2 and the setup of [Agarwal2024ManyShotIL], using 4-shot as a reference and scaling up to 100 examples. On MagpieMath, we observe 71.2 (0-shot), 71.4 (4-shot), 72.8 (10-shot), a peak at 75.2 (15-shot), followed by 71.2 (20-shot), 70.8 (30-shot), 70.0 (50-shot), 70.6 (80-shot), and 70.6 (100-shot). The rise-then-fall trajectory replicates the central empirical regularity documented in prior work. The observations above are consistent with [Agarwal2024ManyShotIL]. In their Section 3.1, Figure 7 (left), Gemini 1.5 Pro on MATH-500 with ground-truth exemplars exhibits the same shape: performance rises as the number of shots increases, peaks at roughly 125 shots, and then declines as the context grows further. Although [Agarwal2024ManyShotIL] truncates at 500 shots, the downward trend is already evident. In our experiments on MagpieMath, when we extend beyond 30–100 examples the decline persists and eventually dips below the 4-shot baseline, noting that the absolute scores differ across datasets and evaluation protocols. Taken together, evaluations spanning Qwen2.5-7B, Qwen2.5-14B, GPT-4o-mini, and previously reported results for Gemini 1.5 Pro all exhibit the same qualitative regularity: performance improves, peaks, and then degrades as the number of in-context examples grows. As model capability increases, the peak tends to occur at a larger shot count and the subsequent degradation is slower, suggesting that stronger long-context handling delays but does not eliminate many-shot brittleness. These results reinforce that long-context capacity is necessary but not sufficient; robust many-shot generalization remains an open challenge.
Appendix D Clarifying “Passive Consumption” (LfI) vs. “Experience-Driven Adaptation” (LfE)
Although both settings can appear as few-shot style prompts, they differ in intent and supervision: LfI (passive consumption) supplies instructional content—worked solutions, step-by-step rationales, or corrections—from a strong instructor, providing pedagogically curated guidance; formally, the context is $\mathcal{C}_{t}=\mathcal{I}_{t}$ and prediction follows $f_{\theta}(x,\mathcal{I}_{t})\!\to\!\hat{y}$. LfE (experience-driven adaptation) instead presents episodic traces—raw input–output pairs, partial trajectories, or outcomes—without didactic restructuring, requiring the model to infer what to abstract or discard; here $\mathcal{C}_{t}=\tau_{t}=\{s_{1},\ldots,s_{k}\}$ and prediction follows $f_{\theta}(x,\tau_{t})\!\to\!\hat{y}$. In our experiments, LfI uses instructor-produced solutions/rationales, whereas LfE uses selected histories from earlier rounds as off-policy trajectories, reflecting guided transfer versus autonomous abstraction from unstructured experience.
Appendix E Rationale for Dataset Choices Across Learning Dimensions
The evaluation follows a two-stage design. Sections 4–6 isolate and validate each learning mechanism under task families that provide the right supervision affordances and minimal cross-factor interference. Section 7 then integrates all three mechanisms in a unified setting via LearnArena, enabling a holistic assessment. A single fixed dataset for all experiments would entangle constructs and reduce interpretability, whereas the staged design improves construct validity by matching task structure to the learning signal being tested. For Learning from Instructor (Section 4), mathematics is used because it naturally supports a tutor–learner paradigm with high-fidelity supervision. Problems admit worked solutions and stepwise rationales, enabling clear, pedagogically oriented instructional signals from a strong instructor (e.g., Qwen2.5-Math-72B) and objective evaluation of whether instruction improves the learner. For Learning from Concept (Section 5), domains are chosen where structured knowledge can be injected and audited. Competitive environments (TextArena) expose explicit rules and strategy templates; logic puzzles (LogicGrid, NaLogic) emphasize formal constraints; and planning/agent tasks (Plan, AlfWorld, ScienceWorld, BabyAI) probe procedural abstractions and generalization. These settings make conceptual cues precise, testable, and meaningfully tied to downstream behavior. For Learning from Experience (Section 6), the focus is on adaptation from trajectories rather than instruction. Two complementary regimes are used: on-policy experience in competitive games, where histories accumulate across rounds and must be exploited under stochastic interaction, and off-policy experience via in-context examples, where pre-collected traces serve as episodic supervision without didactic restructuring. This separation targets the specific cognitive demands of experience integration. Section 7 introduces LearnArena as a unified benchmark that combines instructional feedback, conceptual summaries, and experience selection within shared game environments. The unified evaluation complements the isolated studies by testing whether the three learning mechanisms compose in realistic multi-agent settings, while the earlier sections justify that each mechanism is measurable and interpretable in its most suitable domain.
Appendix F Why Three Perspectives? Scope and Relationship to Self-Critique
[…]